OpenStudy (anonymous):

what is a span , basis and and subspace ? , if C spans B and B spans A . SUCH THAT A , B ,C ARE VECTORS IN A SUBSPACE D ? .

OpenStudy (anonymous):

I will First Clear Your Doubt on BASIS AND SUBSPACE

OpenStudy (anonymous):

1. Do the vectors j1 = (1,0,-1,2) and j2 = (0,1,1,2) form a basis for the space W = {(a,b,c,d) l a - b + c = 0, -2a - 2b + d =0} ?

OpenStudy (anonymous):

answer:1. i substitute each vector in a,b,c,d equation and find the equality, so happen when i substitute in all returns 0, which implies that it does form a basis. Is there a proper working for this, cause i can do this by inspection.

OpenStudy (anonymous):

2 BASIS Find a basis for \[W = { (a,b,c,d) : a - b + 2d = 0 , 3a + c + 3d = 0 }\]

OpenStudy (anonymous):

answer: I was wondering forming a matrix and row reduced the matrix and get a linear dependence equation and plug in some value for c & d to find a & b.

OpenStudy (anonymous):

3. Find the dimension of the following subspaces of R^3

OpenStudy (anonymous):

hint: First, recall the definition of dimension. The dimension of a subspace is the number of linearly independent vectors which span the given subspaces, ie. the number of vectors in the basis of these subspaces. The vectors in any given basis are all linearly independent. Do you see what to do now?

OpenStudy (anonymous):

4. Suppose that a subspace W ⊂ R^3 has a basis { (1,2,3),(1,0,1) }. a)Is { (-1,-2,-3), (3,2,5) } a basis for W? Why?

OpenStudy (anonymous):

answer: HINT: Suppose you are given a basis and another basis which may or may not span the same subspace as that of the first. Informally, if you can express every vector in the first group of vectors as a linear combination of the vectors in the second group, what does that tell you about the corresponding subspace spanned by the latter group with respect to the first?

OpenStudy (anonymous):

if d spans c , where c spans a and , b , i think we can represent every vector in a , b,and c as a linear combination of vectors in d . can we ?

OpenStudy (anonymous):

yes! but my brother you should visit this: http://www.khanacademy.org/math/linear-algebra/v/linear-combinations-and-span

OpenStudy (anonymous):

Man Would you like to be my fan!

OpenStudy (anonymous):

say we choose d in R3 to be (1,2,3,) such that we find any vectors in a ,b,c to represent d

OpenStudy (anonymous):

ya!

OpenStudy (anonymous):

thanks jo , do you know phsyics i am strugling with a question here ?

OpenStudy (anonymous):

ya sure!

OpenStudy (anonymous):

come to this ! http://www.twiddla.com/841565

OpenStudy (anonymous):

http://awwapp.com/draw.html#8198082a

OpenStudy (anonymous):

help